OpenStudy (anonymous):

2(6z-1)-5(z+2)=4(z+1) The Solution is Z= ?

OpenStudy (ranga):

Expand the terms and simplify first: 2(6z-1) = 12z - 2 -5(z+2) = -5z - 10 4(z+1) = 4z + 4 So 12z - 2 -5z - 10 = 4z + 4 7z - 12 = 4z + 4 Subtract 4z from both sides 3z - 12 = 0 + 4 Add 12 to both sides 3z = 16 divide both sides by 3 z = 16/3

OpenStudy (ranga):

Let me put z = 16/3 back in the equation to see if it works...

OpenStudy (ranga):

yes z = 16/3 works.